# GeoBEV: Learning Geometric BEV Representation for Multi-view 3D Object Detection

### v1

@@ Abstract
Bird’s-Eye-View (BEV) representation has emerged as a mainstream paradigm for multi-view 3D object detection, demonstrating impressive perceptual capabilities. However, existing methods overlook the geometric quality of BEV representation, leaving it in a low-resolution state and failing to restore the authentic geometric information of the scene. In this paper, we identify the reasons why previous approaches are constrained by low BEV representation resolution and propose Radial-Cartesian BEV Sampling (RC-Sampling), enabling efficient generation of high-resolution dense BEV representations without the need for complex operators. Additionally, we design a novel In-Box Label to substitute the traditional depth label generated from the LiDAR points. This label reflects the actual geometric structure of objects rather than just their surfaces, injecting real-world geometric information into the BEV representation. Furthermore, in conjunction with the In-Box Label, a Centroid-Aware Inner Loss (CAI Loss) is developed to capture the fine-grained inner geometric structure of objects. Finally, we integrate the aforementioned modules into a novel multi-view 3D object detection framework, dubbed GeoBEV. Extensive experiments on the nuScenes dataset exhibit that GeoBEV achieves state-of-the-art performance, highlighting its effectiveness.
1 Introduction
@@
Multi-view 3D object detection stands as a prominent perception paradigm for cost-effective autonomous driving. Presently, many camera-only detectors (Huang et al. 2021; Huang and Huang 2022a; Li et al. 2023c, 2022, b; Yang et al. 2023) transform image features into Bird’s-Eye-View (BEV) space and directly perform detection on the BEV features, demonstrating competitive performance. This illustrates the substantial advantages of BEV representation in preserving comprehensive scene information, making it more adept for vision-centric autonomous driving perception than isolated image features in perspective space (Park et al. 2021; Wang et al. 2021, 2022b, 2022a).
@@
Figure 1: Comparison between BEV representations. BEVDepth (Li et al. 2023c) is chosen as the baseline. Larger BEV size, RC-Sampling and In-Box Label are gradually applied on the baseline. The boxes represent the ground truth of the scene and brightness in BEV representation reveals the norm of the features. The background is filtered out in (b)-(e) to show the difference in the foreground.
@@
As the cornerstone of BEV-based approaches, the BEV representation embodies both contextual semantic information and depth geometric information. The former is derived from transformed image features, while the latter originates from the correlation between image features and BEV features, calculated using camera parameters. Both types of information are indispensable for precise 3D object detection. However, the geometric quality of BEV representation has never received sufficient attention, and the limitations of low-resolution representations always arise: (1) For LSS-based methods that transform pseudo-points into BEV representation (Xie et al. 2022; Philion and Fidler 2020; Huang et al. 2021; Huang and Huang 2022a), the density imbalance of pseudo-points results in many positions within the BEV representation having vacant features, as depicted in the peripheral areas of Fig. 1(a). Furthermore, the sparsity will further increase along with the resolution of the BEV representation, as shown in Fig. 1(c). (2) For Transformer-based methods that employ cross-attention to retrieve image features (Li et al. 2022; Yang et al. 2023; Jiang et al. 2023; Li et al. 2023a), the elevated resolution of the BEV representation leads to a rapid escalation in computational costs. The lack of explicit depth distribution also limits their ability to restore accurate geometric information. Overall, regardless of the type of method, their respective shortcomings restrict the generation of high-resolution BEV representations, thus compromising the ability to accurately depict objects.
To solve the drawbacks of the existing feature transformation mechanisms, we propose Radial-Cartesian BEV Sampling (RC-Sampling) to efficiently generate dense BEV representation with high resolution. Initially, Radial BEV features are obtained through high-dimensional matrix multiplication between the transposed image features and depth scores. Subsequently, bilinear sampling is employed to retrieve the corresponding Radial BEV features for populating the BEV features in Cartesian coordinates. The sampling procedure mitigates the sparsity of the BEV representation and incurs minimal computational cost even with increased resolution. As shown in Fig. 1(d), applying RC-Sampling effectively avoids the vacant features that appear in Fig. 1(c), enhancing the geometric quality of BEV representations. RC-Sampling outperforms state-of-the-art feature transformation approaches on both precision and efficiency without using complex operators.
@@
Enhancing the accuracy of geometric information is equally crucial for BEV-based approaches. Some methods attempt to supervise predicted depth scores by utilizing the depth values of LiDAR points as depth labels (Reading et al. 2021; Li et al. 2023c, b; Wang et al. 2022c; Zhang et al. 2023a). However, the LiDAR labels only record the depth of object surfaces near the camera, failing to represent the actual geometric structure of objects in the real-world space. We propose In-Box Label that can significantly enhance the geometric quality of BEV representations without adopting extra modules. We first check whether the generated pseudo-points are within the GT boxes to obtain binary labels. These labels, called vanilla In-Box Label, can effectively incentivize the network to assign high depth scores to where the objects are actually located. Nonetheless, they may lead to feature confusion caused by object occlusion or background pseudo-points being wrongly boxed. We ameliorate those issues to enhance its accuracy in reflecting the geometric structure of the scene. In conjunction with the utilization of In-Box Labels, Centroid-Aware Inner Loss is also proposed to capture the fine-grained inner geometric structure of objects. After applying the In-Box Label, the authentic geometric structures of objects are clearly presented as shown in Fig. 1(e), and more precise detection is facilitated.
@@
  • We propose Radial-Cartesian BEV Sampling to conveniently acquire Cartesian BEV features by bilinearly sampling Radial BEV features, which enables the efficient generation of high-resolution dense BEV representations, facilitating the recovery of fine-grained geometric details within the scene.
@@
  • We design a novel In-Box Label to supervise the predicted depth score, which better reflects the actual geometric structure of the object than the LiDAR label, injecting authentic geometric information into the BEV representation.
@@
  • Extensive experiments are conducted on the nuScenes Dataset, and GeoBEV reaches newly state-of-the-art results for multi-view 3D object detection, highlighting its effectiveness.
2 Related Work
2.1 Depth Prediction Based BEV Representation
@@
Due to the inherent limitations of the image modality in providing the indispensable explicit depth information required for 3D object detection, predicting the depth distribution of image elements becomes a natural choice. Early methods like OFT (Roddick, Kendall, and Cipolla 2018) assume that the depth distribution of image elements is uniform and all voxels along the ray starting from the camera share the same features. Lately, LSS (Philion and Fidler 2020) enables networks to adaptively predict depth distributions and weight image features to generate pseudo-points at corresponding depth values, which are then accumulated into BEV representations. BEVDet (Philion and Fidler 2020) employs LSS to construct the detection framework and proposes a data augmentation strategy on BEV features. BEVDet4D (Huang and Huang 2022a) integrates the BEV features from past frames to assist in predicting the velocity of the objects in the current frame.
@@
In order to obtain more accurate depth information, CaDDN (Reading et al. 2021) projects the LiDAR points onto the image to provide supervision on the predicted depth distribution. BEVDepth (Li et al. 2023c) also considers the influence of the camera’s internal and external parameters and further optimizes the depth distribution after the supervised prediction. BEVStereo (Li et al. 2023b) utilizes the structure of multi-view stereo to obtain more reliable depth distributions and performs some optimizations to minimize memory usage. TiG-BEV (Huang et al. 2022) sets several key points to learn the local depth structure of the scene better. SA-BEV (Zhang et al. 2023a) segments the images to get the foreground-only BEV features and improves depth distribution via multi-task learning. BEV-IO (Zhang et al. 2023b) adopts instance occupancy prediction modules as a complement to depth prediction. FB-BEV (Li et al. 2023d) combines forward and backward projection, compensating for the deficiencies in both existing methods. However, these attempts to optimize depth prediction fail to avoid geometrical fallacies. In this paper, we propose In-Box Labels that contain comprehensive geometric information, significantly enhancing the geometric perception of BEV representation.
@@
Figure 2: Overall architecture of GeoBEV. The multi-view images are processed to provide the image features and depth scores. The depth scores are supervised by the In-Box Label that restores authentic geometric structures of objects through the Centroid-Aware Inner Loss. Radial-Cartesian BEV Sampling then conveniently generates dense BEV representation with high resolution.
2.2 Transformer Based BEV Representation
@@
With the help of the attention mechanism, Transformer-based detectors can adaptively retrieve and combine image features to obtain dense BEV representation. BEVFormer (Li et al. 2022) uses deformable attention to find the corresponding image features, and additionally fuses the BEV representation from previous frames. BEVFormerV2 (Yang et al. 2023) introduces a detection head in perspective view to make the image features that will be retrieved more suitable for 3D detection. PolarFormer (Jiang et al. 2023) argues that polar coordinates are more competent for ego car perception and designs a polar detection head to deal with the irregular polar grids. DFA3D (Li et al. 2023a) utilizes the explicit depth distribution when doing the cross attention from image to BEV and simplifies the 3D Transformer into the 2D Transformer equivalently.
@@
Several Transformer-based detectors regard the objects as queries and do not generate explicit BEV representation. DETR3D (Wang et al. 2022b) follows DETR (Carion et al. 2020; Zhu et al. 2020) series detectors and interacts object queries with multi-view image features to complete the object information for detection. PETR (Liu et al. 2022) embeds 3D position information into the image features, making the object queries easier to obtain spatial information. PETRv2 (Liu et al. 2023a) extends PETR for temporal modeling and adds map queries for other perception tasks. StreamPETR (Wang et al. 2023) propagates long-term historical information and achieves comparable accuracy to classical LiDAR based detectors. Sparse4D (Lin et al. 2022) assigns multiple 4D key points to aggregate multi-view/scale/timestamp image features. Sparse4Dv2 (Lin et al. 2023) uses the recurrent method to realize the transmission of temporal information, avoiding multiple sampling to improve efficiency. Nevertheless, the existing Transformer-based BEV representation methods involve complex computation processes and high latency, making it difficult to generalize to high-resolution BEV feature situations. In this paper, we propose Radial-Cartesian BEV Sampling to efficiently generate high-resolution dense BEV representations, facilitating the recovery of fine-grained geometric details within the scene.
3 Method
3.1 Overall Architecture
@@
The overall architecture of our proposed GeoBEV is shown in Fig. 2. Firstly, the multi-view images are processed to provide the image features that will be transformed into BEV features and their corresponding depth scores. Then the In-Box Label is created and utilized to supervise the depth scores to restore the actual distribution of the objects in BEV space. The Centroid-Aware Inner Loss is adopted during the depth supervision to let the model learn the inner structure of the objects. Finally, Radial-Cartesian BEV Sampling is proposed to conveniently generate dense BEV representation with high resolution, which effectively saves the fine-grained geometric information of the scene.
@@
Figure 3: The illustration of Radial-Cartesian BEV Sampling. After high-dimensional matrix multiplication between the transposed image features and depth scores, the $H$ dimension is squeezed and the radial BEV features can be directly obtained.
3.2 Radial-Cartesian BEV Sampling
Increasing the resolution of BEV features can significantly enhance its detail and make it easier for perception modules to extract fine-grained geometric information. However, the drawbacks of current methods limit the resolution of BEV representations. For LSS-based methods (Li et al. 2023c, b; Zhang et al. 2023a; Li et al. 2023d), the density imbalance of pseudo-points makes the signal strength of BEV features decays from the center to the periphery, and some places even have the vacant features. The sparsity will be further increased for BEV features in high resolution. FB-BEV (Li et al. 2023d) applies backward projection to fill these vacant features but relies on precisely predicting RoI from the sparse BEV features. Transformer-based methods(Li et al. 2022; Yang et al. 2023; Jiang et al. 2023; Qin et al. 2023) can create dense BEV features, but the computational cost increases rapidly when the resolution of BEV features increases.
Here, we propose Radial-Cartesian BEV Sampling (RC-Sampling) to conveniently generate dense BEV features with high resolution. The radial BEV features are first obtained by high-dimensional matrix multiplication between transposed image features and the predicted depth scores as shown in Fig. 3. Regardless of batch size, the image features can be denoted as $\mathbb{I}\in\mathbb{R}^{C\times H\times W}$, where $C,H,W$ represent the channel number, height and weight of the image features. The depth scores can be denoted as $\mathbb{D}\in\mathbb{R}^{D\times H\times W}$, where $D$ represents the number of discrete depth values. To get the radial BEV features, the $H$ dimension should be squeezed, so that the image features and depth scores are transposed to $\hat{\mathbb{I}}\in\mathbb{R}^{W\times C\times H}$ and $\hat{\mathbb{D}}\in\mathbb{R}^{W\times H\times D}$ respectively. The high-dimensional matrix multiplication is then applied to the last two dimensions of $\hat{\mathbb{I}}$ and $\hat{\mathbb{D}}$ and a new matrix is created by:
$\hat{\mathbb{B}}_{R}=\hat{\mathbb{I}}\hat{\mathbb{D}},$ (1)
where $\hat{\mathbb{B}}_{R}\in\mathbb{R}^{W\times C\times D}$. After transposing $\hat{\mathbb{B}}_{R}$ into $\mathbb{B}_{R}\in\mathbb{R}^{C\times D\times W}$, the radial BEV features are obtained.
@@ 3.2 Radial-Cartesian BEV Sampling
The $B_{R}$ needs to be transformed into Cartesian coordinates for subsequent detection. We pre-define the coordinates of Cartesian BEV features $\mathbb{B}_{C}\in\mathbb{R}^{C\times X\times Y}$, where $X,Y$ denote the required resolution, and simply project them on the $\mathbb{B}_{R}$. The bilinear sampling is utilized to retrieve the corresponding features, which can be represented by :
$\mathbb{B}_{C}(x,y)={\rm BilinearSample}(\mathbb{B}_{R},{\rm Project}(x,y)),$ (2)
@@
where ${\rm Project}(x,y)$ denotes the coordinates of the projected point $(x,y)$ on $\mathbb{B}_{R}$. Using bilinear sampling instead of pooling the sparse pseudo-points guarantees that each position in the BEV representation has valid information. The sampled $\mathbb{B}_{C}$ can represent the geometric information of the scenes well because the explicit depth distribution has already been incorporated into the $\mathbb{B}_{R}$. When $\mathbb{B}_{C}$ with high resolution are used for perception, the $\mathbb{I}$ and $\mathbb{D}$ with high resolution are also required to provide fine-grained geometric information of the scene. In practice, several convolutions are utilized to increase $H$ and $W$ of $\mathbb{I}$ and $\mathbb{D}$.
@@
Compared with other feature transformation methods, RC-Sampling does not require the generation of memory-expensive 4D frustum features (Reading et al. 2021; Philion and Fidler 2020; Huang et al. 2021), the utilization of deployment-unfriendly custom operators (Li et al. 2023c; Huang and Huang 2022b; Liu et al. 2023b) or the computation-expensive cross-attention mechanism (Li et al. 2022; Yang et al. 2023; Li et al. 2023a). It is entirely the combination of the most basic official operators. RC-Sampling not only shows the advantage of generating high-resolution BEV representation, but also outperforms the state-of-the-art feature transformation methods, such as BEVPoolv2 (Huang and Huang 2022b) and DFA3D (Li et al. 2023a), on both precision and efficiency.
(a) LiDAR Label
(b) Vanilla Label
(c) Occlusion
(d) Background
(e) Behind Surface
(f) CAI Loss
@@
Figure 4: Illustration of the associated design of In-Box Label. $H$, $W$ and $D$ represent the height, width and channel number of depth scores. The boxes are the GT boxes. The red squares and black dots denote the positive and negative classes of the In-Box Label. The triangles are the LiDAR points. The crosses in blue represent the pseudo-points that are not supervised. The deeper color in the last figure means higher loss weight.
3.3 In-Box Label
@@
Current LSS-based detectors (Li et al. 2023c, b, d; Zhang et al. 2023a) utilize the depth of LiDAR points as the label to supervise the depth score of each pseudo-point and thus attach the geometric information of the scene to the BEV representation. However, as shown in Fig. 4(a), the LiDAR label only records the depth of the object surfaces closer to the camera, instead of the actual geometric structure of the objects. The lack of objects’ complete geometric information hinders the subsequent BEV encoder and detection head from precisely recognizing their size and orientation. Besides, the LiDAR points do not exactly match pseudo-points in space and will introduce errors that can not be eliminated. To overcome the drawbacks of the LiDAR label, we propose the In-Box Label that can be easily obtained from the 3D coordinate of pseudo-points and the GT boxes.
@@
Denote the 3D coordinate of a pseudo-point generated from image features as $p\in\mathbb{R}^{3}$ and the space within a GT box as $B$, then the vanilla In-Box Label of $p$ can be formulated as:

### v2

@@ Abstract
Bird’s-Eye-View (BEV) representation has emerged as a mainstream paradigm for multi-view 3D object detection, demonstrating impressive perceptual capabilities. However, existing methods overlook the geometric quality of BEV representation, leaving it in a low-resolution state and failing to restore the authentic geometric information of the scene. In this paper, we identify the drawbacks of previous approaches that limit the geometric quality of BEV representation and propose Radial-Cartesian BEV Sampling (RC-Sampling), which outperforms other feature transformation methods in efficiently generating high-resolution dense BEV representation to restore fine-grained geometric information. Additionally, we design a novel In-Box Label to substitute the traditional depth label generated from the LiDAR points. This label reflects the actual geometric structure of objects rather than just their surfaces, injecting real-world geometric information into the BEV representation. In conjunction with the In-Box Label, Centroid-Aware Inner Loss (CAI Loss) is developed to capture the inner geometric structure of objects. Finally, we integrate the aforementioned modules into a novel multi-view 3D object detector, dubbed GeoBEV, which achieves a state-of-the-art result of 66.2% NDS on the nuScenes test set. The code is available at https://github.com/mengtan00/GeoBEV.git.
Introduction
@@
Multi-view 3D object detection stands as a prominent perception paradigm for cost-effective autonomous driving. Presently, many camera-only detectors (Huang et al. 2021; Huang and Huang 2022a; Li et al. 2023c, 2022, b; Yang et al. 2023) transform image features into Bird’s-Eye-View (BEV) space and directly perform detection on the BEV features, demonstrating competitive performance. This illustrates the substantial advantages of BEV representation in preserving comprehensive scene information, making it more adept for vision-centric autonomous driving perception than isolated image features in perspective space (Park et al. 2021; Wang et al. 2021, 2022a, 2022b).
(a) Baseline
(b) Larger BEV size
(c) Applying RC-Sampling
(d) Applying In-Box Label
@@
Figure 1: Comparison between BEV representations. BEVDepth is chosen as the baseline. Larger BEV size, RC-Sampling and In-Box Label are added in turn. The boxes represent the ground truth of the scene and brightness reveals the norm of the features. The background is filtered out to show the difference in the foreground.
@@
As the cornerstone of BEV-based approaches, the BEV representation embodies both contextual semantic information and depth geometric information. The former is derived from image features, while the latter originates from the correlation between image features and BEV features. Both types of information are indispensable for precise 3D object detection. However, the geometric quality of BEV representation has never received sufficient attention, and the limitation of low-resolution representation always arises. LSS-based methods  (Xie et al. 2022; Philion and Fidler 2020; Huang et al. 2021; Huang and Huang 2022a) pool pseudo-points into BEV representation, leaving the positions without pseudo-points to have vacant features. The sparsity will further increase along with the BEV resolution as shown in Fig. 1(b). For Transformer-based methods (Li et al. 2022; Yang et al. 2023; Jiang et al. 2023; Li et al. 2023a) that employ cross-attention to retrieve image features, the elevated BEV resolution leads to a rapid escalation in computational costs. The lack of explicit depth distribution also limits their ability to restore accurate geometric information. Some methods (Harley et al. 2023; Peng et al. 2023) simply sample the image features to obtain voxel features, which are then squeezed into BEV representation. However, this kind of feature transformation, which we call Voxel-Sampling, requires a large number of sampling operations and produces huge intermediate features, giving it no advantage over LSS-based and Transformer-based methods.
To solve the drawbacks of the existing feature transformation mechanisms, we propose Radial-Cartesian BEV Sampling (RC-Sampling) to generate dense BEV representation with high resolution efficiently. Initially, we create Radial BEV features correlated to each camera view by extending the depth dimension and squeezing the height dimension of image features. We prove that this step can be achieved by simple matrix transposition and multiplication without creating huge intermediate voxel features. Subsequently, bilinear sampling is employed to retrieve the corresponding Radial BEV features for populating the BEV features in Cartesian coordinates. The number of sampling operations between different BEV features is far less than the sampling operations used to create voxel features. RC-Sampling creates the BEV representation of the same quality as Voxel-Sampling while reducing more than 90% time cost and memory cost according to our experiments. RC-Sampling is also faster than the most efficient LSS approach like BEVPoolv2 (Huang and Huang 2022b), which relies on custom operator acceleration, and thoroughly solves the problem of the feature vacancy as shown in Fig. 1(c).
@@
Truthfully representing the real spatial distribution of the objects is as important as increasing the BEV resolution. Some methods supervise predicted depth scores by utilizing the depth values of LiDAR points as depth labels (Reading et al. 2021; Li et al. 2023c, b; Wang et al. 2022c; Zhang et al. 2023a; Jiang et al. 2025). However, the LiDAR labels only record the depth of object surfaces that face the ego car, failing to represent the actual geometric structure of objects in real-world space. We propose In-Box Label to offer more competent supervision. We first check whether the generated pseudo-points are within the GT boxes and obtain binary labels. These labels, called Vanilla In-Box Label, can effectively incentivize the network to assign high depth scores to where the objects are actually located. Nonetheless, they may lead to feature confusion caused by object occlusion or wrongly boxed background pseudo-points. We ameliorate those issues to enhance its accuracy in reflecting the geometric structure of the scene. In conjunction with the utilization of In-Box Labels, Centroid-Aware Inner Loss (CAI Loss) is also proposed to capture the fine-grained inner geometric structure of objects. After applying the In-Box Label, the authentic geometric structures of objects are clearly presented as shown in Fig. 1(d), and more precise detection is facilitated. It is noteworthy that both In-Box labels and CAI Loss do not introduce extra parameters.
@@
  • We propose Radial-Cartesian BEV Sampling to conveniently acquire Cartesian BEV features by bilinearly sampling Radial BEV features, which enables the efficient generation of high-resolution dense BEV representation, facilitating the recovery of fine-grained geometric details within the scene.
@@
  • We design the novel In-Box Label, cooperating with Centroid-Aware Inner Loss, to supervise the depth scores, which better reflects the actual geometric structure of the object than the LiDAR label and inject authentic geometric information into the BEV representation.
@@
  • Extensive experiments are conducted on the nuScenes Dataset, and GeoBEV reaches newly state-of-the-art results of 66.2% NDS for multi-view 3D object detection, highlighting its effectiveness.
Related Work
Depth Prediction Based BEV Representation
@@
Due to the limitation of the camera in capturing the depth required for 3D object detection, predicting the depth distribution of image elements becomes a natural choice. Early methods like OFT (Roddick, Kendall, and Cipolla 2018) assume that the depth distribution of image elements is uniform and all voxels along the ray share the same features. Lately, LSS (Philion and Fidler 2020) enables adaptive depth prediction and weights image features to generate pseudo-points at corresponding depth values, which are pooled into BEV features. BEVDet (Huang et al. 2021) employs LSS to construct the detection framework and applies data augmentation in the BEV space. BEVDet4D (Huang and Huang 2022a) merges the BEV features from past frames to help predict the objects’ velocity.
@@
In order to obtain more accurate depth information, CaDDN (Reading et al. 2021) projects the LiDAR points onto the image to supervise the predicted depth distribution. BEVDepth (Li et al. 2023c) considers the camera’s internal and external parameters and further optimizes the depth distribution after the supervision. BEVStereo (Li et al. 2023b) introduces multi-view stereo to obtain more reliable depth distributions and performs some optimizations to minimize memory usage. TiG-BEV (Huang et al. 2022) sets key points to learn the local depth structure of the scene. SA-BEV (Zhang et al. 2023a) segments the images to get the foreground-only BEV features and improves depth distribution via multi-task learning. BEV-IO (Zhang et al. 2023b) adopts instance occupancy prediction modules as a complement to depth prediction. FB-BEV (Li et al. 2023d) adds a backward process to fill a part of vacant features. BEVNext (Li et al. 2024) adopts CRF to modulate the estimated depth. However, these attempts fail to record the actual object structure due to the limitation of LiDAR points.
@@
Figure 2: Overall architecture of GeoBEV. The multi-view images are processed into the image features and depth scores. The depth scores are supervised by the In-Box Label that records authentic geometric structures of objects through the Centroid-Aware Inner Loss. Radial-Cartesian BEV Sampling then efficiently generates dense BEV representation with high resolution.
Transformer Based BEV Representation
@@
With the attention mechanism, Transformer-based detectors can adaptively retrieve image features to obtain dense BEV representation. BEVFormer (Li et al. 2022) uses deformable attention to transform image features into BEV space and fuse the past BEV representation. BEVFormerV2 (Yang et al. 2023) introduces a detection head in perspective view to make the image features more suitable for 3D detection. PolarFormer (Jiang et al. 2023) generates the BEV representation in polar coordinates, which are more competent for ego car perception. DFA3D (Li et al. 2023a) utilizes the explicit depth distribution in cross attention and simplifies the 3D Transformer into the 2D Transformer equivalently.
@@
Several Transformer-based detectors regard the objects as queries to save the large amount of computation required to generate the explicit BEV representation. DETR3D (Wang et al. 2022b) follows DETR series detectors (Carion et al. 2020; Zhu et al. 2020) and interacts object queries with multi-view image features. PETR (Liu et al. 2022) embeds 3D position into the image features, supplementing spatial information to the object queries. PETRv2 (Liu et al. 2023) extends PETR for temporal modeling and adds map queries for other perception tasks. StreamPETR (Wang et al. 2023) propagates long-term historical information. Sparse4D (Lin et al. 2022) assigns multiple 4D key points to aggregate multi-view/scale/timestamp image features. Sparse4Dv2 (Lin et al. 2023) uses the recurrent method to transmit the temporal information. RayDN (Liu et al. 2024) constructs positive and negative examples along the camera rays to learn depth-aware features. Nevertheless, the omission of explicit BEV representation causes geometric information loss, limiting their precision upper bound.
Method
Overall Architecture
@@
The overall architecture of our proposed GeoBEV is shown in Fig. 2. Firstly, the multi-view images are processed by the image backbone and DepthNet to provide the image features and depth scores. Then the In-Box Label is created and utilized to supervise the depth scores to restore the actual distribution of the objects in BEV space. Centroid-Aware Inner Loss is adopted to let the model learn the inner structure of the objects. Finally, Radial-Cartesian BEV Sampling generates dense BEV representation with high resolution, outperforming current feature transformation approaches in both efficiency and effectiveness.
@@
Figure 3: The illustration of Radial-Cartesian BEV Sampling. After high-dimensional matrix multiplication between the transposed image features and depth scores, the $H$ dimension is squeezed to obtain Radial BEV features.
Radial-Cartesian BEV Sampling
The drawbacks of current methods limit the resolution of BEV representation, failing to restore the fine-grained geometric information of the scene. For LSS-based methods, the density imbalance of pseudo-points leads to feature vacancy in BEV representation, which will deteriorate further as the BEV resolution increases. FB-BEV (Li et al. 2023d) applies backward projection to fill these vacant features but relies on imprecise RoI predicted from the sparse BEV features. The cross-attention between the image and BEV space in Transformer-based methods guarantees the density of BEV features, but the computational cost increases rapidly along with the BEV resolution. Some methods  (Peng et al. 2023; Harley et al. 2023) sample image features following strict projection relations to obtain voxel features, which are summarized along the height into dense BEV features. However, the large number of sampling operations and the huge intermediate features lessen their practicability.
Here, we propose Radial-Cartesian BEV Sampling (RC-Sampling) to generate dense BEV features with high resolution conveniently. Firstly, the Radial BEV features corresponding to each image view are created by extending the depth dimension and summarizing the height dimension of the image features. They are so-called because their elements appear radially distributed after being projected into the 3D space as shown in Fig 3. We denote image features as $\mathbb{I}\in\mathbb{R}^{\scriptscriptstyle C\times H\times W}$ and depth scores as $\mathbb{D}\in\mathbb{R}^{\scriptscriptstyle D\times H\times W}$, where $C,D,H,W$ represent the channel, depth, height and width dimension respectively. The elements in Radial BEV features $\mathbb{B}^{R}\in\mathbb{R}^{\scriptscriptstyle C\times D\times W}$ can be represented by:
$\gamma_{cdw}=\sum_{h\in H}\alpha_{chw}\beta_{dhw},$ (1)
where $\alpha,\beta,\gamma$ are the elements of $\mathbb{I},\mathbb{D},\mathbb{B}^{R}$, and $c,d,h,w$ are the indexes of $C,D,H,W$. The general approach is to create 4D frustum features $\mathbb{F}\in\mathbb{R}^{\scriptscriptstyle C\times D\times H\times W}$ and summarize along the $H$ dimension. By contrast, RC-Sampling implements this process more efficiently.
Omitting the common dimension $W$, the Equation 1 can be simplified as:
$\gamma_{cd}=\sum_{h\in H}\alpha_{ch}\beta_{dh}\Rightarrow\mathbb{B}^{R}_{w}=\mathbb{I}_{w}\mathbb{D}^{\top}_{w},$ (2)
where $\mathbb{B}^{R}_{w}\in\mathbb{R}^{\scriptscriptstyle C\times D},\mathbb{I}_{w}\in\mathbb{R}^{\scriptscriptstyle C\times H},\mathbb{D}_{w}\in\mathbb{R}^{\scriptscriptstyle D\times H}$ are the slices of $\mathbb{I},\mathbb{D},\mathbb{B}^{R}$ at $w$. After considering dimension $W$, $\mathbb{B}^{R}$ can be directly created by:
$\mathbb{B}^{R}=[(\mathbb{I}\rightarrow\mathbb{R}^{\scriptscriptstyle W\times C\times H})\otimes(\mathbb{D}\rightarrow\mathbb{R}^{\scriptscriptstyle W\times H\times D})]\rightarrow\mathbb{R}^{\scriptscriptstyle C\times D\times W},$ (3)
where $\rightarrow$ and $\otimes$ denote the transposition and multiplication of the high-dimension matrix. As shown in Fig. 3, additional computation and memory required by frustum features $\mathbb{F}$ are saved for equally creating $\mathbb{B}^{R}$.
@@
The $\mathbb{B}^{R}$ needs to be transformed into Cartesian coordinates for subsequent detection. We pre-define the coordinates of Cartesian BEV features $\mathbb{B}^{C}\in\mathbb{R}^{\scriptscriptstyle C\times X\times Y}$, where $X,Y$ denote the required BEV resolution, and project them on the $\mathbb{B}^{R}$. The bilinear sampling is utilized to retrieve the corresponding features, which can be represented by :
$\mathbb{B}^{C}(x,y)={\rm BilinearSample}(\mathbb{B}^{R},{\rm Project}(x,y)),$ (4)
@@
where ${\rm Project}(x,y)$ denotes the coordinates of the projected point $(x,y)$ on $\mathbb{B}^{R}$. Bilinearly sampling $\mathbb{B}^{R}$ instead of pooling the sparse pseudo-points guarantees that each position in $\mathbb{B}^{C}$ has valid features. It also saves more than 90% time cost and memory cost required by Voxel-Sampling while creating BEV representation with equal geometric quality. The efficiency and quality advantages can be maintained when the BEV resolution is increased. Along with larger $\mathbb{B}^{C}$, $\mathbb{I}$ and $\mathbb{D}$ are also enlarged by lightweight convolution to provide fine-grained information of the scene.
@@
Compared with other feature transformation methods, RC-Sampling does not require the generation of memory-expensive 3D intermediate features, the utilization of deployment-unfriendly custom operators or the computation-expensive cross-attention mechanism, highlighting its usability. Experiment results illustrate that RC-Sampling outperforms the state-of-the-art feature transformation methods, such as BEVPoolv2 and DFA3D, on both precision and efficiency.
(a) Occlusion Between Objects
(b) Background in Box
(c) Behind Background Surface
(d) Centroid-Aware Inner Loss
@@
Figure 4: Illustration of the associated design of In-Box Label. $H,W,D$ represent the height, width and depth dimensions. The boxes are the GT boxes. The red squares and black dots denote the positive and negative points of the In-Box Label. The blue crosses are the points that are not supervised. The deeper color in (d) means higher loss weight.
In-Box Label
@@
LiDAR points have been used to supervise the depth score of each pseudo-points, effectively attaching the geometric information to the BEV representation. However, the LiDAR label only records the depth of the object surfaces facing the ego car, instead of the actual geometric structure of the objects. The lack of objects’ complete geometric information hinders the subsequent BEV encoder and detection head from precisely recognizing their size and orientation. To overcome the drawbacks of the LiDAR label, we propose the In-Box Label that can be easily obtained from the 3D coordinate of pseudo-points and the GT boxes.
@@
Denote the 3D coordinate of a pseudo-point generated from image features as $p\in\mathbb{R}^{3}$ and the space within a GT box as $B$, the Vanilla In-Box Label can be formulated as: